# Tin-Based Chalcogenide Perovskites: A Promising Lead-Free Alternative for Stable and High-Performance Photovoltaics

### v1

@@ III.2 Electronic Properties:
After examining the stability, the electronic properties, such as band structure and partial density of states (PDOS) of the chalcogenide perovskites ASnX3 (A = Ca, Sr, Ba; X = S, Se) are computed to gain deep insights for designing the photoelectric devices. At first, the electronic band structure calculations are performed using the semi-local GGA-PBE xc functional with and without including the spin-orbit coupling (SOC) for these CPs. It is found that GGA-PBE xc functional is unable to predict the correct bandgaps due to the self-interaction error of the electrons (see Table 2), and SOC does not have any impact on the bandgap (see Table S6 of the SM), which is expected for chalcogenide perovskites[22, 14]. After that, we employed the hybrid HSE06 xc functional and many-body perturbation theory (MBPT) based GW (G0W0@PBE) method to calculate the bandgaps more accurately. The HSE06 and G0W0@PBE calculated band structures of these compounds are depicted in Figure 2 and Figure S3, respectively. Our results revealed that Ca- and Sr-based compounds have direct bandgaps at (0, 0, 0) point, while Ba-based CPs exhibit indirect bandgaps. In the case of BaSnS3 and BaSnSe3, the valence band maxima (VBM) are situated at the S (0.5, 0.5, 0) and U (0.5, 0, 0.5) points of the Brillouin zone (BZ), respectively, while the conduction band minima (CBM) are found in between and X (0.5, 0, 0) point for both cases (see Figure 2). Also, both of them exhibit the lowest direct bandgap in between and X point of the BZ. The HSE06 as well as G0W0@PBE estimated bandgaps of these CPs are listed in Table 2, which closely match with earlier theoretical and available experimental findings. For example, the G0W0@PBE bandgap of CaSnS3 (1.44 eV) closely aligns with its experimental bandgap of 1.72 eV[32]. The discrepancy in bandgap values could be attributed to the temperature, as the experimental bandgap was measured at 500° C. The HSE06 and G0W0@PBE computed bandgap of these CPs lie in the range of 0.701.45 eV and 0.791.50 eV, respectively. Thus, the bandgaps of the investigated CPs are within the ideal range for photovoltaic applications, aligning well with those of most lead-based HPs (1.503.13 eV)[23, 24, 28], as well as Zr- and Hf-based CPs (1.692.46 eV)[14, 22, 15] (for details, see Table S13 of the SM). This suggests their high suitability for photovoltaic applications.
@@
Figure S2 shows the partial density of states (PDOS) and the total density of states (TDOS) for all the studied compounds obtained using the HSE06 xc functional. In ASnS3 (A = Ca, Sr, Ba) CPs, S- orbitals mostly contribute to the VBM, while the hybridization of S- and Sn- orbitals dominates the CBM. Similarly, for ASnSe3 perovskites, the VBM is primarily contributed by Se- orbitals, whereas CBM mainly consists of hybridized Se- and Sn- orbitals. The hybridization of Se- and Sn- orbitals in ASnSe3 is stronger than the hybridization of S- and Sn- orbitals in ASnS3, which is mainly responsible for lowering the bandgap in case of ASnSe3 as compared to their sulfur counterparts ASnS3.
@@
Notably, perovskites are highly recognized for their ability to transport materials. Therefore, in addition to their band structures, we also calculate the corresponding electron () and hole () effective masses by utilizing the fitted E−k dispersion band diagram of G0W0@PBE band structures (Figure S3), employing the formula, (for details, see the SM). From Table 3, it is observed that < 1, < 1.5 for all the cases, indicating high carrier mobility and thus better charge carrier transport. Since BaSnX3 (X = S, Se) exhibits an indirect bandgap, the effective masses for these compounds are calculated at both of their indirect and direct band edges. Also, Table 3 suggests that Ca- and Sr-based CPs have a high potential for exhibiting ambipolar characteristics.
@@ III.3 Optical Properties:
The above study shows that the electronic properties of these CPs can be well-described by the HSE06 functional. Nevertheless, this functional is known to predict the optical features of these systems with less accuracy. Thus, we performed MBPT-based GW-BSE calculations to calculate the optical properties. In essence, GW calculations compute the fundamental bandgap, which is thought to be more accurate and comparable to photoelectron spectroscopy (PES) and inverse photoelectron spectroscopy (IPES)[40, 41], while BSE calculations predict the optical bandgap similar to experimental optical absorption spectroscopy[42, 43].
To evaluate the optical responses of the ASnX3 CPs, BSE calculations are performed on top of the single-shot GW(G0W0)@PBE, which explicitly considers the electron-hole interaction. The real [Re ()] and imaginary [Im ()] part of the frequency-dependent electronic dielectric function calculated using BSE@G0W0@ PBE are shown in Figure3. It is discovered that the absorption onset and the first peak position () gradually red shift from sulfide (S) to selenide (Se) containing CPs akin to the drop of quasiparticle (QP) bandgap of them (see Table 2). For example, the first peak position for CaSnS3 is observed at 1.23 eV, and it shifts to 0.74 eV for CaSnSe3. The value of for these ASnX3 CPs falls within the range of 0.74 to 1.26 eV and shifts with the change of the A atom, similar to the variation observed in their QP bandgap (see Table 4). The electronic dielectric constants (), which are obtained from the real part of the dielectric function, also found to be increased from S to Se containing CPs indicating low charge carrier recombination rate and improved optoelectronic efficiency for the latter[70] (see Table 4). For instance, for CaSnS3 is 7.47, and it increases to 15.15 for CaSnSe3. Overall, these results reveal that CaSnS3, SrSnS3, BaSnS3, and BaSnSe3 could be the best choice for photovoltaic applications.
@@ III.4 Excitonic Properties:
In optoelectronic materials, exciton generation plays a major role in the charge separation properties and for this reason, excitonic parameters, such as exciton binding energy () and exciton lifetime (), are crucial descriptors in their applications. The energy required to dissociate an exciton into a single electron () and hole () pair is known as the exciton binding energy. Theoretically, is computed as the difference of the QP band gap (direct G0W0@PBE bandgap, ) and the optical band gap (BSE@G0W0@PBE peak position, )[71, 72, 56, 73], which are tabulated in Table 4. It should be noted that the electronic contribution of the dielectric screening is dominant over the ionic one if is significantly greater than the longitudinal optical phonon energy (). In that scenario, one can neglect the ionic contribution and consequently, remains unchanged[74, 75]. From Table 4 and Table 6, it is found that for ASnX3 CPs; therefore, ionic screening to the dielectric function can be disregarded. The values obtained from standard first-principles GW-BSE calculations are found in the range of 0.050.24 eV, which are comparable to that of Zr- and Hf-based CPs (0.020.26 eV)[14, 22, 15] but a little higher than those observed in conventional lead-based HPs (0.010.10 eV)[23, 24, 25] (for details, see Table S13 of the SM), which ensure the potential of these materials for the optoelectronic applications. Here, CaSnSe3 exhibits the lowest value, attributed to its higher electronic dielectric constant (), which reduces the Coulomb interaction between the electron and hole.
@@
In addition, we have validated the using the Wannier-Mott method[22, 56] through the Eq. 1. Here, lies in between optical () and static () dielectric constant. The optical or electronic dielectric constants () at the zero frequency limit are obtained using BSE method. Subsequently, the DFPT calculations are performed to compute the ionic contribution to the dielectric function (), and therefore, the static dielectric constant is calculated as, = ( + ). In our study, the upper () and lower () bounds of exciton binding energy are estimated based on the and , respectively (see Table 4). The upper bound values are smaller but align more closely with the calculated using the GW-BSE method compared to the lower bound values; however, the overall trend remains consistent. This suggests that in chalcogenide perovskites, the electronic contribution to dielectric screening is more prominent than the ionic contribution.
Furthermore, additional investigations into the ionic (phonon) contribution to the exciton binding energy () are conducted for the examined CPs. This is because the standard BSE approach within an ab-initio framework captures only static screening from electrons to calculate the exciton binding energy. However, it’s noteworthy that dynamic electron-electron interactions or electron-phonon coupling may play a crucial role in certain materials, particularly those with significant electron-phonon interactions or where phonons significantly influence optoelectronic properties. Filip and collaborators[57] have recently broadened the discussion to incorporate phonon screening into the analysis of exciton binding energy using equation 2. They achieved this by considering four specific material parameters: reduced effective mass, static and optical dielectric constants, and the frequency of the longitudinal optical phonon mode (), while assuming isotropic and parabolic electronic band dispersion. Table 5 indicates that phonon screening leads to a reduction in the ranging from 5.58% to 17.87%, with the modified values ranging from 0.04 to 0.23 eV. While phonon screening does decrease the exciton binding energy, the reduction is not particularly substantial, except in the case of CaSnSe3. This also suggests that in the majority of chalcogenide perovskites, the electronic contribution to dielectric screening outweighs the ionic (or phonon) contribution.
@@
Several excitonic parameters, including excitonic temperature (), exciton radius (), and probability of wave function () for pair at zero charge separation are also computed using the above quantities (, , and ) to have a definitive estimation of the excitonic properties (for details, see the SM). The inverse of can be used to qualitatively characterize the exciton lifetime (), which is listed in Table 4 (for details, see the SM). Consequently, the values for the investigated CPs are in the order CaSnSe3 SrSnSe3 CaSnS3 BaSnSe3 SrSnS3 BaSnS3. A longer exciton lifetime corresponds to a lower carrier recombination rate, which enhances the quantum yield and conversion efficiency. Overall, these properties significantly enhance the efficiency of ASnX3 CPs, making them promising for potential optoelectronic applications.
@@ III.5 Polaronic Properties:
In polar semiconductors, such as halide perovskite and its derivatives, the scattering mechanism near room temperature is dominant due to the interaction between charge carriers and the macroscopic electric field produced by longitudinal optical phonon (LO)[59]. This interaction strongly influences charge carrier mobility of the system and is anticipated to be same for the materials of our interest[14, 22]. Therefore, the Fröhlich mesoscopic model[59, 26, 76] is used to describe this interaction and defined by the dimensionless Fröhlich parameter using Eq. 3. The calculated values of related to Fröhlich interaction for electrons and holes are given in Table 6 and Table S11, respectively. Strong electron (hole)-phonon coupling is indicated by , while often suggests weak coupling[26]. Our results show that the value of for ASnX3 CPs lies in the range of 0.642.92, suggesting weak to intermediate electron (hole)-phonon coupling for the investigated systems. Here, CaSnSe3 exhibits smaller electron-phonon coupling, while BaSnS3 demonstrates larger electron-phonon coupling. Weak electron-phonon coupling can be attributed to factors such as a lower electron effective mass and a higher electronic dielectric constant. Conversely, strong coupling arises from the opposite factors. In this context, the specific free volume of these CPs are also evaluated to obtain a qualitative knowledge of the strength of electron-phonon coupling (for details, see the SM).
@@
The other parameters for the polarons, i.e., polaron mass () and polaron mobility (), are also estimated using Eq. 5 and Eq. 6, respectively. These parameter values are listed in Table 6 and Table S11 for electrons and holes, respectively. One can confirm the enhanced carrier-lattice interactions by higher values, and for this case, the charge carrier mobility decreases than the non-polar or less polar compounds. For example, Ca- and Sr-based compounds exhibit weaker carrier (electron)-lattice interactions compared to Ba-based compounds, resulting in higher mobility for Ca- and Sr-based materials. Also, CaSnSe3 CP exhibit highest polaron mobility for electron which is obvious due to weak electron-phonon coupling ( = 0.64). Overall, These Sn-based materials are shown to have ambipolar properties and much improved polaron mobility for electrons (21.33416.02 cm2V-1s-1) and holes (7.02260.69 cm2V-1s-1, see Table S11 of the SM) than the conventional lead-based HPs (57290 cm2V-1s-1 for electrons and 97230 cm2V-1s-1 for holes, respectively)[26, 27] as well as Zr- and Hf-based CPs (6.8477.59 cm2V-1s-1 for electrons and 3.76100.49 cm2V-1s-1 for holes, respectively)[14, 22, 15].

### v2

@@ III.2 Electronic Properties:
After examining the stability, the electronic properties, such as band structure and partial density of states (PDOS) of the chalcogenide perovskites ASnX3 (A = Ca, Sr, Ba; X = S, Se) are computed to gain deep insights for designing the photoelectric devices. At first, the electronic band structure calculations are performed using the semi-local GGA-PBE xc functional with and without including the spin-orbit coupling (SOC) for these CPs. It is found that GGA-PBE xc functional is unable to predict the correct bandgaps due to the self-interaction error of the electrons (see Table 2), and SOC does not have any impact on the bandgap (see Table S6 of the Supplemental Material), which is expected for chalcogenide perovskites[22, 14]. After that, we employed the hybrid HSE06 xc functional and many-body perturbation theory (MBPT) based GW (G0W0@PBE) method to calculate the bandgaps more accurately. The HSE06 and G0W0@PBE calculated band structures of these compounds are depicted in Figure 2 and Figure S3, respectively. Our results revealed that Ca- and Sr-based compounds have direct bandgaps at (0, 0, 0) point, while Ba-based CPs exhibit indirect bandgaps. In the case of BaSnS3 and BaSnSe3, the valence band maxima (VBM) are situated at the S (0.5, 0.5, 0) and U (0.5, 0, 0.5) points of the Brillouin zone (BZ), respectively, while the conduction band minima (CBM) are found in between and X (0.5, 0, 0) point for both cases (see Figure 2). Also, both of them exhibit the lowest direct bandgap in between and X point of the BZ. The HSE06 as well as G0W0@PBE estimated bandgaps of these CPs are listed in Table 2, which closely match with earlier theoretical findings. It is well known that the G0W0@PBE method provides highly accurate bandgap values for CPs, yielding results that closely match experimental values[14, 22, 15]. However, the G0W0@PBE bandgap of CaSnS3 (1.44 eV) is in reasonable agreement with its experimental bandgap of 1.72 eV[32], showing a difference of 0.28 eV. The underestimation of the GW bandgap may partly arise from structural discrepancies (c-lattice constant), as GW calculations are sensitive to input geometries, and deviations from experimental structures can affect electronic properties. Additionally, the experimental bandgap of CaSnS3 was measured at 500°C (773.15 K), while theoretical calculations are performed at 0 K, and this temperature difference could also contribute to the discrepancy. Furthermore, incorporating excitonic effects into the calculations could potentially lead to a larger difference between the theoretical (optical) and experimental bandgaps. The HSE06 and G0W0@PBE computed bandgap of these CPs lie in the range of 0.701.45 eV and 0.791.50 eV, respectively. Thus, the bandgaps of the investigated post-transition metal Sn-based CPs fall within the optimal range for photovoltaic applications, aligning closely with those of conventional APbI3 (A = Cs, CH3NH3) HPs (1.501.85 eV)[23, 28] and being lower than the bandgaps of transition metal based CPs (1.692.46 eV)[14, 22, 15] (for details, see Table S14 of the Supplemental Material). This suggests their high suitability for photovoltaic applications.
@@
Figure S4 shows the partial density of states (PDOS) and the total density of states (TDOS) for all the studied compounds obtained using the HSE06 xc functional. In ASnS3 (A = Ca, Sr, Ba) CPs, S- orbitals mainly contribute to the VBM, while the hybridization (sp3-type) of S- and Sn- orbitals dominates the CBM. Similarly, for ASnSe3 perovskites, the VBM is primarily contributed by Se- orbitals, whereas CBM mainly consists of hybridized Se- and Sn- orbitals. The hybridization of Se- and Sn- orbitals in ASnSe3 is stronger than the hybridization of S- and Sn- orbitals in ASnS3, which is mainly responsible for lowering the bandgap in the case of ASnSe3 as compared to their sulfur counterparts ASnS3. On the other hand, in the case of transition metal-based CPs, the bonding (sp3d2) is a mix of bonding (from the covalent overlap of orbitals) and bonding (to a lesser extent, from -orbital interactions), leading to strong covalent interactions. At the same time, a more significant difference in the electronegativity of Zr/Hf and S/Se adds some ionic characteristics (see Figure S5), which eventually results in more localized electronic states around the atom and, thus, larger bandgaps than Sn-based CPs[14, 22, 15].
@@
Notably, perovskites are highly recognized materials for efficiently transporting charge carriers (electrons and holes). Therefore, in addition to their band structures, we also calculate the corresponding electron () and hole () effective masses by utilizing the fitted E−k dispersion band diagram of G0W0@PBE band structures (Figure S3), employing the formula, (for details, see the Supplemental Material). From Table 3, it is observed that < 1, < 1.5 for all the cases, indicating high carrier mobility and thus better charge carrier transport. Since BaSnX3 (X = S, Se) exhibits an indirect bandgap, the effective masses for these compounds are calculated at both their indirect and direct band edges. Also, Table 3 suggests that Ca- and Sr-based CPs have a high potential for exhibiting ambipolar characteristics.
@@ III.3 Optical Properties:
In addition to the electronic bandgaps, we also explore the optical bandgaps to gain a holistic understanding of the excitonic effects in perovskite materials. These effects, arising from the interaction of electrons and holes, are pivotal for determining the efficiency of light-matter interactions in optoelectronic applications. The above study shows that the electronic properties of these CPs can be well-described by the HSE06 functional. Nevertheless, this functional is known to predict the optical features of these systems with less accuracy. Thus, we performed MBPT-based GW-BSE calculations to calculate the optical properties. In essence, GW calculations compute the fundamental bandgap, which is thought to be more accurate and comparable to photoelectron spectroscopy (PES) and inverse photoelectron spectroscopy (IPES)[40, 41]. In contrast, BSE calculations predict the optical bandgap similar to experimental optical absorption spectroscopy[42, 43].
To evaluate the optical responses of the ASnX3 CPs, BSE calculations are performed on top of the single-shot GW(G0W0)@PBE, which explicitly considers the electron-hole interaction. The real [Re ()] and imaginary [Im ()] part of the frequency-dependent electronic dielectric function calculated using BSE@G0W0@ PBE are shown in Figure3. It is discovered that the absorption onset and the first peak position or optical bandgap () gradually red shift from sulfide (S) to selenide (Se) containing CPs akin to the drop of quasiparticle (QP) bandgap of them (see Table 2). For example, the optical bandgap for CaSnS3 is observed at 1.23 eV, and it shifts to 0.74 eV for CaSnSe3. After incorporating the excitonic effect in CaSnS3, it is evident that the difference between the theoretical (optical) bandgap and the experimental bandgap increases by 0.49 eV. However, the underlying reason for this difference remains consistent with the previously discussed influence of structural geometry and temperature. The value of for these ASnX3 CPs falls within the range of 0.74 to 1.26 eV and shifts with the change of the A atom, similar to the variation observed in their QP bandgap (see Table 4). Our calculations reveal that the optical bandgaps of these CPs are consistently smaller than their electronic bandgaps due to significant excitonic effects. This study primarily focuses on the bright excitonic states, however, dark excitons can significantly influence the optical properties of semiconductors, particularly in systems with low symmetry or strong spin-orbit coupling. Dark excitons, which are optically inactive due to spin-forbidden transitions or momentum mismatch, are typically found below the bright excitons in the absorption spectrum. Although they do not couple directly to light, their presence can affect the overall photophysical behavior, such as photoluminescence and carrier dynamics. Nevertheless, an analysis of the BSE eigenvalues revealed no evidence of dark excitons (optically inactive states) below the first bright exciton for the investigated materials.
The electronic dielectric constants (), which are obtained from the real part of the dielectric function, are also found to be increased from S to Se containing CPs, indicating low charge carrier recombination rate and improved optoelectronic efficiency for the latter[70] (see Table 4). For instance, for CaSnS3 is 7.47, and it increases to 15.15 for CaSnSe3. Overall, these results reveal that CaSnS3, SrSnS3, BaSnS3, and BaSnSe3 could be the best choice for photovoltaic applications.
@@ III.4 Excitonic Properties:
In optoelectronic materials, exciton generation plays a significant role in the charge separation properties. For this reason, excitonic parameters, such as exciton binding energy () and exciton lifetime (), are crucial descriptors in their applications. The energy required to dissociate an exciton into a single electron () and hole () pair is known as the exciton binding energy. The GW-calculated electronic gaps are complemented by BSE-computed optical gaps to account for excitonic binding energies. This combined approach allows us to predict material properties with high accuracy, ensuring their relevance for practical device design. Theoretically, is computed as the difference of the QP bandgap (direct G0W0@PBE bandgap, ) and the optical bandgap (BSE@G0W0@PBE peak position, )[71, 72, 56, 73], which are tabulated in Table 4. It should be noted that the electronic contribution of the dielectric screening is dominant over the ionic one if is significantly greater than the longitudinal optical phonon energy (). In that scenario, one can neglect the ionic contribution, and consequently, remains unchanged[74, 75]. From Table 4 and Table 6, it is found that for ASnX3 CPs; therefore, ionic screening to the dielectric function can be disregarded. The values obtained from standard first-principles GW-BSE calculations are found in the range of 0.050.24 eV, which are comparable to that of transition metal-based traditional CPs (0.020.26 eV)[14, 22, 15] but a little higher than those observed in conventional lead-based HPs (0.010.10 eV)[23, 24, 25] (for details, see Table S14 of the Supplemental Material). The values obtained from our GW-BSE calculations indicate relatively strong excitonic effects in these materials. While this range may appear too high for direct photovoltaic applications, it is worth noting that these values represent intrinsic properties of isolated bulk materials. In practical device architectures, factors such as dielectric screening from surrounding layers, nanostructuring, or heterostructures can significantly reduce the effective , potentially bringing it within the desired range for photovoltaic applications. Furthermore, these materials may be more suited for other optoelectronic applications, such as light-emitting diodes, photodetectors, or excitonic solar cells, where stronger excitonic effects are advantageous. Here, CaSnSe3 exhibits the lowest value, attributed to its higher electronic dielectric constant (), which reduces the Coulomb interaction between the electron and hole.
@@
In addition, we have validated the using the Wannier-Mott method[22, 56] through the Eq. 1. This method helps establish a qualitative understanding of excitonic properties and their dependence on effective masses and dielectric screening, serving as a sanity check for trends observed in the ab initio BSE calculations. The Wannier-Mott model is traditionally more applicable to direct bandgap materials. However, for indirect bandgap materials in our study, the effective masses and, subsequently, the reduced masses are calculated at their direct band edges (for details, see table 3). Here, lies in between optical () and static () dielectric constant. The optical or electronic dielectric constants () at the zero frequency limit are obtained using the BSE method. Subsequently, the DFPT calculations are performed to compute the ionic contribution to the dielectric function (), and therefore, the static dielectric constant is calculated as, = ( + ). In our study, the upper () and lower () bounds of exciton binding energy are estimated based on the and , respectively (see Table 4). The upper bound values are smaller but align more closely with the calculated using the GW-BSE method than the lower bound values; however, the overall trend remains consistent. This suggests that in chalcogenide perovskites, the electronic contribution to dielectric screening is more prominent than the ionic contribution.
Furthermore, additional investigations into the ionic (phonon) contribution to the exciton binding energy () are conducted for the examined CPs. This is because the standard BSE approach within an ab initio framework captures only static screening from electrons to calculate the exciton binding energy. However, it’s noteworthy that dynamic electron-electron interactions or electron-phonon coupling may play a crucial role in certain materials, particularly those with significant electron-phonon interactions or where phonons significantly influence optoelectronic properties. Filip and collaborators[57] have recently broadened the discussion to incorporate phonon screening into the analysis of exciton binding energy using equation 2. They achieved this by considering four specific material parameters: reduced effective mass, static and optical dielectric constants, and the frequency of the longitudinal optical phonon mode (), while assuming isotropic and parabolic electronic band dispersion. Table 5 indicates that phonon screening leads to a reduction in the ranging from 5.58% to 17.87%, with the modified values ranging from 0.04 to 0.23 eV. While phonon screening does decrease the exciton binding energy, the reduction is not particularly substantial, except in the case of CaSnSe3. This also suggests that in most chalcogenide perovskites, the electronic contribution to dielectric screening outweighs the ionic (or phonon) contribution.
@@
Several excitonic parameters, including excitonic temperature (), exciton radius (), and probability of wave function () for pair at zero separation are also computed using the above quantities (, , and ) to have a definitive estimation of the excitonic properties (for details, see the Supplemental Material). Based on the analysis of exciton radius (), it is observed that in all investigated materials, the electron-hole pairs are distributed over multiple lattice constants, supporting the applicability of the Wannier-Mott model. The inverse of can be used to qualitatively characterize the exciton lifetime (), which is listed in Table 4 (for details, see the Supplemental Material). Consequently, the values for the investigated CPs are in the order CaSnSe3 SrSnSe3 CaSnS3 BaSnSe3 SrSnS3 BaSnS3. For efficient photo-conversion in solar cells, the exciton lifetime () should be sufficiently high to extract photo-generated charge carriers before recombination occurs. A longer exciton lifetime corresponds to a lower carrier recombination rate, which enhances the quantum yield and conversion efficiency. Overall, these properties significantly enhance the efficiency of ASnX3 CPs, making them promising for potential optoelectronic applications.
@@ III.5 Polaronic Properties:
In polar semiconductors, such as halide perovskite and its derivatives, the scattering mechanism near room temperature is dominant due to the interaction between charge carriers and the macroscopic electric field produced by longitudinal optical phonon (LO)[59]. This interaction strongly influences charge carrier mobility of the system and is anticipated to be the same for the materials of our interest[14, 22]. Therefore, the Fröhlich mesoscopic model[59, 26, 76] is used to describe this interaction and defined by the dimensionless Fröhlich parameter using Eq. 3. The calculated values of related to Fröhlich interaction for electrons and holes are given in Table 6 and Table S12, respectively. Strong electron (hole)-phonon coupling is indicated by , while often suggests weak coupling[26]. Our results show that the value of for ASnX3 CPs lies in the range of 0.642.92, suggesting weak to intermediate electron (hole)-phonon coupling for the investigated systems. Additionally, it has been observed that Sn-based CPs exhibit less pronounced polaronic effects compared to transition metal-based CPs[14, 22, 15]. This weaker electron-phonon coupling can be attributed to a lower electron effective mass and a higher electronic dielectric constant, while stronger coupling in transition metal-based CPs results from the opposite factors. Here, CaSnSe3 exhibits smaller electron-phonon coupling, while BaSnS3 demonstrates larger electron-phonon coupling for the same reasons. In this context, the specific free volume of these CPs is also evaluated to obtain a qualitative knowledge of the strength of electron-phonon coupling (for details, see the Supplemental Material).
@@
The other parameters for the polarons, i.e., polaron mass () and polaron mobility (), are also estimated using Eq. 5 and Eq. 6, respectively. These parameter values are listed in Table 6 and Table S12 for electrons and holes, respectively. One can confirm the enhanced carrier-lattice interactions by higher values, and for this case, the charge carrier mobility decreases than the non-polar or less polar compounds. For example, Ca- and Sr-based compounds exhibit weaker carrier (electron)-lattice interactions compared to Ba-based compounds, resulting in higher mobility for Ca- and Sr-based materials. Also, CaSnSe3 CP exhibits the highest polaron mobility for electrons, which is obvious due to weak electron-phonon coupling ( = 0.64). Overall, these post-transition metal Sn-based CPs are shown to have ambipolar properties and much-improved polaron mobility for electrons (21.33416.02 cm2V-1s-1) and holes (7.02260.69 cm2V-1s-1, see Table S12 of the Supplemental Material) than the conventional lead-based HPs (57290 cm2V-1s-1 for electrons and 97230 cm2V-1s-1 for holes, respectively)[26, 27] as well as transition metal (Zr and Hf) based traditional CPs (6.8477.59 cm2V-1s-1 for electrons and 3.76100.49 cm2V-1s-1 for holes, respectively)[14, 22, 15]. The higher polaron mobility in Sn-based CPs is attributed to less prominent polaronic effects than conventional lead-based HPs and transition metal-based CPs.